OpenStudy (anonymous):

The base of a triangle is ten feet longer than the height. The area of the triangle is 100 square feet. Find the height and base of the triangle

OpenStudy (zpupster):

A=1/2bh b=h+10 100 = (1/2)(h+10)(h)